OpenStudy (anonymous):

I'LL GIVE YOU TONS OF MEDALS Given the linear equation 2y + 6 = 5x, find the y-intercept.

OpenStudy (anonymous):

pls someone help i cry

OpenStudy (anonymous):

ill pay in rela money

OpenStudy (kuoministers):

hmm? rela?

OpenStudy (anonymous):

i meant real you it was a typo step off bully

OpenStudy (goten77):

the y intercept is when x=0

OpenStudy (kuoministers):

to find y intercept you let x = 0

OpenStudy (anonymous):

ok hey goten i love yugioh

OpenStudy (anonymous):

also it isnt zero

OpenStudy (goten77):

2y+6=5(0) 2y+6=0 2y=-6 y=-3

OpenStudy (kuoministers):

first step is 2y + 6 = 5 * 0